OpenStudy (anonymous):

How fast must a ball be rolled along the surface of a 70-cm high table so that when it rolls off the edge it will strike the floor at the same distance (70cm) from the point directly below the edge of the table?

OpenStudy (anonymous):

Range = velocity * time of descent. in your question, range is 70 cm and time of descent \[\sqrt{2h / g}\]

OpenStudy (anonymous):

solve the equation.

OpenStudy (anonymous):

@chand do u mean range is \[\sqrt{2h/g}\]

OpenStudy (anonymous):

no. it is the time of descent. range = velocity * time of descent

OpenStudy (mos1635):

 x= Vo * t(disent) x= Vo * sqr(2h/g)